add or subtract as indicated 
  _     ___
2Ö5 + 5Ö125

Answer by AnlytcPhil(1116)

add or subtract as indicated 
  _     ___
2Ö5 + 5Ö125

You can't break 5 down because it's prime, but you can 
break down 125 into primes

125 = 5·25 = 5·5·5, so substitute 5·5·5 for 125
  _     _____
2Ö5 + 5Ö5·5·5

Since it's a square root we group like factors into 
twos. (If it were a cube root you would group into 
threes)

  _     ________
2Ö5 + 5Ö(5·5)·5
 
Since (5·5) is the same as 5², so we can substitute
5² for (5·5) 
        ____
2V5 + 5Ö5²·5

Now we can take individual square roots of the factors
under the second radical
  _     __  _
2Ö5 + 5Ö5²·Ö5

Taking the square root of '5 squared' just takes 
away the square and the radical, and you just have 5:
  _        _
2Ö5 + 5·5·Ö5
  _      _ 
2Ö5 + 25Ö5
                            _
Now you can factor out the Ö5
 _
Ö5(2 + 25)

2 plus 25 is 27
 _
Ö5(27)

Write the 27 first
   _
27Ö5

Edwin